# Towards Generalizable PDE Dynamics Forecasting via Physics-Guided Invariant Learning

### v1

@@ Appendix A LLM Usage
Appendix B Related Work
B.1 Spatiotemporal PDE Dynamics Forecasting
@@
B.2 Invariant Learning for OOD Generalization
@@
B.3 Physical Invariance Learning
@@
B.4 Practical value of Zero-shot OOD Dynamics Forecasting
@@
Appendix C Multi-Environment Dataset Details
@@
Appendix D More Analysis on Mixture of Operator Expert Architecture
D.1 Disparate MoE usage in iMOOE and LFMs
@@
D.2 Detailed explanations on the effect of the number of operator experts
@@
Appendix E Additional Results
E.1 Ablation Study
E.1.1 Effect of Frequency-Enriched Loss in Eq. 8
We investigate the benefits of the proposed frequency enrichment loss $\mathcal{L}_{freq}$ for PDE invariance learning. We utilize DR trajectories to validate the improved forecasting generalizability induced by additional regularization on high-frequency representations. Diffusion and reaction coefficients can dictate the distribution of frequency patterns in DR state evolutions. Apart from nMESE and fRMSE metrics, we also present the forecasting errors within different frequency bands in Table 6. ”Low”, ”Mid”, ”High” denote non-overlapped ranges of wavenumber $\xi$: $\xi_{\text{low}}\in[0,4]$, $\xi_{\text{mid}}\in[5,12]$, $\xi_{\text{high}}\in[13,\xi_{max}]$. After equipping PDE learning with frequency augmentation, the ID/OOD nMSE can drop by $24.38\%$ and $25.00\%$, and ID/OOD fRMSE can decrease by $10.20\%$ and $12.32\%$. Notably, improving high-frequency feature learning can also enhance the OOD accuracy on both low-frequency and mid-frequency patterns. Such ID/OOD promotion verifies the necessity of the proposed frequency-enriched objective, which can mitigate the spectral bias of neural operators and help to capture the complete PDE invariance from the spectral domain.
@@ E.1 Ablation Study
E.1.2 Effect of Pre-calculated Derivative Selection
@@
iii) Actual learned mask vectors $\mathbf{m}$. When learning on DR data, we feed second-order derivatives $u_{xx},v_{xx},u_{yy},v_{yy},u_{xy},v_{xy}]$ to two experts in iMOOE, and their actual learned mask is $\mathbf{m}_{1}=[0,0,1,1,1,0]$, $\mathbf{m}_{2}=[1,1,1,0,0,1]$. As there are many operator splitting methods for DR equation (e.g. dividing into diffusion and reaction terms is just one of them), iMOOE can learns a suitable splitting way via learning operator invariance from limited data.
@@
E.1.3 Effect of the Choice of Fusion Network
@@
E.1.4 Effect of Step-Wise Environment Partition
As mentioned in Section 3.2, dividing the training environments based on autoregressive time steps can further boost the outcomes of PDE invariance learning. We adopt two fluid dynamics datasets, namely DR and BG to verify this step-wise partition method. During the fluid evolution, the state variations on two consecutive time steps are quite distinct, but the physical transition law between these two steps remain invariant. Therefore, we can regard each autoregressive step as a unique context. It can enhance the diversity of training environments and lead to more robust PDE invariance representations. In Table 9, we exhibit the benefits brought by such step-wise environment division. We can see that adding this environment partition method to original invariance training can result in an average increase of $10.41\%$ and $4.09\%$ on ID and OOD generalization performance for these two fluid dynamics systems.
Table 9: Ablation results of step-wise environment division on NS and BG data.
Methods NS BG
@@ E.1.4 Effect of Step-Wise Environment Partition
w/o step-wise division 7.11e-2 1.50e-2 3.41e-1 5.49e-2 1.48e-3 1.17e-3 1.11e-2 3.94e-3
w/ step-wise division 6.49e-2 1.38e-2 3.12e-1 5.36e-2 1.20e-3 1.10e-3 1.08e-2 3.83e-3
E.1.5 Effect of Linear Loss Scheduling
@@
E.2 Applicability to Diverse Dynamical System Forecasting
@@
E.2.1 Application to ODE-governed Dynamics Forecasting
We verify iMOOE’s zero-shot OOD capability on neural ODE systems, as prior meta-learning-based methods such as CoDA (Kirchmeyer et al., 2022) and GEPS (Kassaï Koupaï et al., 2024) are also extended to ODE-governed dynamics forecasting. We conduct simulation on a typical ODE system called damped and driven pendulum equation, as shown in Appendix B.1 of (Kassaï Koupaï et al., 2024). We utilize past $10$-step states to forecast the pendulum motion angle of future $41$ steps. The time horizon of collected trajectories is $[0,25]$ and $N_{t}=51$. We construct 16 ID training domains and 8 OOD test domains by randomly drawing four ODE parameters from the ID/OOD ranges given in Table 12. Identical to CoDA and GEPS, a 4-layer MLP network with 64 hidden dimension is taken as the backbone for operator experts of iMOOE. The pre-calculated spatial derivatives and mask diversity loss are discarded since they are unnecessary for ODE simulation. We report OOD forecasting performance of zero-shot iMOOE and few-shot CoDA, GEPS in Table 12. iMOOE can achieve 10.4% and 12.66% increase on nMSE and fRMSE versus GEPS. This may stem from their difference on discovering physical invariance. Specifically, iMOOE explicitly prescribes the two-level invariance principle and directly captures it via the proposed physics-informed invariant learning. While hypernetwork-based meta-learning methods like CoDA, GEPS estimate such invariance by implicitly operating in the network parameter space without any physical guidance.
@@ E.2.1 Application to ODE-governed Dynamics Forecasting
E.2.2 Application to 3D Fluid Dynamics Forecasting
Apart from the typical 2D PDE dynamics, we also demonstrate iMOOE’s zero-shot OOD performance on 3D PDE systems. We employ the 3D compressible Navier-Stokes equation in PDEBench (Takamoto et al., 2022) and construct ID and OOD scenarios using the same method in Appendix C. Specifically, for shear and bulk viscosity coefficients, we still randomly draw their values from the ID parameter range [1e5,1e3][1e-5,1e-3] and OOD range $5e-6,8e-6]\cup[1.2e-3,2e-3]$. The Mach number is kept as 1.0. The number of ID training and OOD test domains as well as their data volume are also identical to setups in Appendix C. The size of 3D spatial domain is $32\times 32\times 32$, and the past 10-step velocity field sequences are provided to forecast the future 11-step states. FNO3d is utilized as the backbone of operator experts for iMOOE3d. We present the 3D OOD dynamics forecasting results in Table 14. We observe that iMOOE can attain 17.36% and 39.26% increase on nMSE and fRMSE compared to previous 3D neural operators. Such results further validate the effectiveness of proposed physics-guided PDE invariance learning to more complex 3D PDE dynamics.
E.2.3 Application to Real-world Time Series Prediction
@@
E.3 More Results on Sensitivity Analysis
By hyperparameter tuning, we empirically find the fixed setting $\lambda_{pred}=1.0,\lambda_{inv}=0.001,\lambda_{freq}=0.1,\lambda_{mask}=0.001$ can perform well on both simulated and real-world physical dynamics data. As the effect of $\mathcal{L}_{mask}$ has been discussed in Appendix E.1.2, we investigate iMOOE’s sensitivity to different $\lambda_{inv}$ and $\lambda_{freq}$ values. OOD results in Table 15 reflect that we should assign moderate values to $\lambda_{inv}$ and $\lambda_{freq}$ for satisfactory outcomes.
@@ E.3 More Results on Sensitivity Analysis
E.4 Further comparison with Meta-learning-based Methods
@@
E.5 Empirical Upper Bound of Zero-shot OOD Performance
@@
E.6 More Results on ID-OOD Correlations in Fig. 1(c)
@@
E.7 Analysis on Training Data Properties
@@
For the second question, we keep the total number of training samples at 1,024 and alter the number of training environments from 4 to 512. The training data quantity in each environment is equal. We depict the distribution of ID and OOD results of each test sample in Fig. 5(b), 6(b). Overall, with diverse training domains, i.e. when the number of training environments is up to 32, ID/OOD results of each test trajectory can disperse more compactly. In other words, the variance across test domains is smaller, and the average ID/OOD fRMSE is much lower. This reveals that when data budget is limited, better data diversity can avoid overfitting to limited training domains, and aid to find the fundamental PDE invariance principle by equalizing the risks across more diverse training environments. This is coherent with the key claim in foundational invariant learning literature (Arjovsky et al., 2019): with a sufficiently large number of diverse training environments, invariant risk minimization would elicit the invariant predictor.
@@ E.9 Implementation Details
E.10 Visualization on OOD Forecasting Results of iMOOE
In Fig. 7 to 13, we visualize the forecasting outcomes of iMOOE on representative OOD physical environments of the five PDE dynamical systems.

### v2

@@ Appendix A LLM Usage
Appendix B Rigorous justification for two-level PDE invariance
Firstly, in the field of PDE learning, the commonly utilized assumption is that many specialized PDE systems demonstrate compositional structures and contains a group of distinct physical processes, as claimed in Section 4.1 of (McCabe et al., 2024). For example, Navier-Stokes systems are composed of convection, pressure and viscosity, wave systems are composed of Hamiltonian flow and dissipation. This PDE compositionality assumption can be formulated as (Xu et al., 2025): $\partial_{t}\mathbf{u}(\mathbf{x})=F\left(\sigma_{1}(\mathbf{u},\mathbf{x}),...,\sigma_{i}(\mathbf{u},\mathbf{x}),...,\sigma_{K}(\mathbf{u},\mathbf{x})\right)$. $F(\cdot)$ is a nonlinear function, which represents the compositional relationship among a group of basic operators $\{\sigma_{i}(\mathbf{u},\mathbf{x})\}_{i=1}^{K}$. Such compositional operator structure guides us to design the two-level PDE invariance principle. The proposed iMOOE leverages a fusion network to approximate $F(\cdot)$ and a set of neural operator experts to approximate $\{\sigma_{i}(\mathbf{u},\mathbf{x})\}_{i=1}^{K}$.
Secondly, as observed in classical PDE discovery literature such as SINDy (Brunton et al., 2016) and PDE-FIND (Rudy et al., 2017), PDE systems of interests can be naturally represented in a few spatial operators. These PDE discovery methods resort to SVD-based linear combination of a few terms in the fixed operator library. This indeed motivates us to design the operator invariance principle, and also justifies the compositional structure of operator learning.
Thirdly, we justify the composition of neural operators (i.e. iMOOE’s architecture design) can approximate arbitrary continuous operators according to the universal approximation theorem proved in (Kovachki et al., 2023). As illustrated in Section 9.3 and Figure 16 in (Kovachki et al., 2023), neural operators that can approximate continuous operators should possess three modules: i) Mapping input to the first hidden representation by a local operator. ii) Mapping each hidden representation to next by a non-local integral kernel operator. iii) Mapping the last hidden representation to output by a local operator. In practical implementation, i) and iii) can be realized by a local neural network like MLP, while ii) should be realized by a global neural network like transformer or FNO layers. The proposed composition of neural operators realizes i) and ii) by concatenating the last hidden representations of multiple neural operators, and realizes iii) by the fusion network. Hence, iMOOE architecture design can also satisfy the universal approximation theorem of operators.
Appendix C Related Work
C.1 Spatiotemporal PDE Dynamics Forecasting
@@
C.2 Invariant Learning for OOD Generalization
@@
C.3 Physical Invariance Learning
@@
C.4 Practical value of Zero-shot OOD Dynamics Forecasting
@@
Appendix D Multi-Environment Dataset Details
@@
Appendix E More Analysis on Mixture of Operator Expert Architecture
E.1 Disparate MoE usage in iMOOE and LFMs
@@
E.2 Detailed explanations on the effect of the number of operator experts
@@
Appendix F Additional Results
F.1 Ablation Study
F.1.1 Effect of Frequency-Enriched Loss in Eq. 8
We investigate the benefits of the proposed frequency enrichment loss $\mathcal{L}_{freq}$ for PDE invariance learning. We utilize the simulated DR data and real-word SST, SSE data to validate the improved forecasting generalizability induced by additional regularization on high-frequency representations. Diffusion and reaction coefficients can dictate the distribution of frequency patterns in DR evolutions. SSE contains high-frequency short waves which could be caused by the nonlinear surface features like wave breaking fronts, sharp crests and bound harmonics. SST contains high-frequency modes due to the ocean advection and vertical processes such as upwelling. Apart from nMESE and fRMSE metrics, we also present the forecasting errors within different frequency bands in Table 7, 8. ”Low”, ”Mid”, ”High” denote non-overlapped ranges of wavenumber $\xi$: $\xi_{\text{low}}\in[0,4]$, $\xi_{\text{mid}}\in[5,12]$, $\xi_{\text{high}}\in[13,\xi_{max}]$. When equipped with high-frequency augmentation, the ID/OOD nMSE can drop by $24.38\%$ and $25.00\%$, and ID/OOD fRMSE can decrease by $10.20\%$ and $12.32\%$ on DR data. It also improves nMSE by $9.22\%$ and $8.98\%$, fRMSE by $7.03\%$ and $7.33\%$ for SST, SSE data. Notably, improving high-frequency feature learning can also enhance the OOD accuracy on both low-frequency and mid-frequency patterns. Such ID/OOD promotion verifies the necessity of the proposed frequency-enriched objective, which can mitigate the spectral bias of neural operators and help to capture the complete PDE invariance from the spectral domain.
@@ F.1 Ablation Study
Table 8: Ablation results of frequency enrichment loss on real-word ocean dynamics data.
Methods SST SSE
nMSE fRMSE nMSE fRMSE
Low Mid High Total Low Mid High Total
w/o $\mathcal{L}_{freq}$ 5.64e-1 1.29e-1 3.55e-2 9.24e-3 3.70e-2 1.67e-2 3.03e-3 2.98e-3 3.01e-3 3.00e-3
w/ $\mathcal{L}_{freq}$ 5.12e-1 1.24e-1 3.06e-2 8.47e-3 3.44e-2 1.52e-2 2.10e-3 2.74e-3 2.94e-3 2.78e-3
F.1.2 Effect of Pre-calculated Derivative Selection
@@
iii) Actual learned mask vectors $\mathbf{m}$. When learning on DR data, we feed second-order derivatives $u_{xx},v_{xx},u_{yy},v_{yy},u_{xy},v_{xy}]$ to two experts in iMOOE, and their actual learned mask is $\mathbf{m}_{1}=[0,0,1,1,1,0]$, $\mathbf{m}_{2}=[1,1,1,0,0,1]$. As there are many operator splitting methods for DR equation (e.g. dividing into diffusion and reaction terms is just one of them), iMOOE can learn a suitable splitting way via learning operator invariance from limited data.
@@
F.1.3 Effect of the Choice of Fusion Network
@@
F.1.4 Effect of Two Environment Partition Methods
As mentioned in Section 3.2, dividing the training environments based on autoregressive time steps can further boost the outcomes of PDE invariance learning. We adopt two fluid dynamics datasets to verify the benefit of this step-wise partition method in addition to common parameter-based division. During the fluid evolution, state variations on two consecutive time steps are quite distinct, but the physics transition law between these two steps remain invariant. Therefore, we can regard each autoregressive step as a unique context. In Table 11, we present the effect of two environment partition methods. We can see that combining two partition methods together can realize the best ID/OOD performance, since it can enhance the diversity of training environments and improve the robustness of learned PDE invariance representations. Besides, parameter-based partition performs moderately better than step-wise partition, as different physical parameters can lead to more distinct PDE trajectories, such as the Reynold number in NS is a decisive factor to distinguish laminar or turbulent flow.
Table 11: Effect of two environment partition methods on fluid forecasting.
Partition methods NS BG
@@ F.1.4 Effect of Two Environment Partition Methods
Only parameters 7.11e-2 1.50e-2 3.41e-1 5.49e-2 1.48e-3 1.17e-3 1.11e-2 3.94e-3
Only time steps 7.39e-2 1.52e-2 3.58e-1 5.52e-2 1.57e-3 1.19e-3 1.18e-2 4.01e-3
Parameters+time steps 6.49e-2 1.38e-2 3.12e-1 5.36e-2 1.20e-3 1.10e-3 1.08e-2 3.83e-3
F.1.5 Effect of Linear Loss Scheduling
@@
F.2 Applicability to Diverse Dynamical System Forecasting
@@
F.2.1 Application to ODE-governed Dynamics Forecasting
We verify iMOOE’s zero-shot OOD capability on neural ODE systems, as prior meta-learning-based methods such as CoDA (Kirchmeyer et al., 2022) and GEPS (Kassaï Koupaï et al., 2024) are also extended to ODE-governed dynamics forecasting. We conduct simulation on a typical ODE system called damped and driven pendulum equation, as shown in Appendix B.1 of (Kassaï Koupaï et al., 2024). We utilize past $10$-step states to forecast the pendulum motion angle of future $41$ steps. The time horizon of collected trajectories is $[0,25]$ and $N_{t}=51$. We construct 16 ID training domains and 8 OOD test domains by randomly drawing four ODE parameters from the ID/OOD ranges given in Table 14. Identical to CoDA and GEPS, a 4-layer MLP network with 64 hidden dimension is taken as the backbone for operator experts of iMOOE. The pre-calculated spatial derivatives and mask diversity loss are discarded since they are unnecessary for ODE simulation. We report OOD forecasting performance of zero-shot iMOOE and few-shot CoDA, GEPS in Table 14. iMOOE can achieve 10.4% and 12.66% decrease on nMSE and fRMSE versus GEPS. This may stem from their difference on discovering physical invariance. Specifically, iMOOE explicitly prescribes the two-level invariance principle and directly captures it via the proposed physics-informed invariant learning. While hypernetwork-based meta-learning methods like CoDA, GEPS estimate such invariance by implicitly operating in the network parameter space without any physical guidance.
@@ F.2.1 Application to ODE-governed Dynamics Forecasting
F.2.2 Application to 3D Fluid Dynamics Forecasting
Apart from the typical 2D PDE dynamics, we also demonstrate iMOOE’s zero-shot OOD performance on 3D PDE systems. We employ the 3D compressible Navier-Stokes equation in PDEBench (Takamoto et al., 2022) and construct ID and OOD scenarios using the same method in Appendix D. Specifically, for shear and bulk viscosity coefficients, we still randomly draw their values from the ID parameter range [1e5,1e3][1e-5,1e-3] and OOD range $5e-6,8e-6]\cup[1.2e-3,2e-3]$. The Mach number is kept as 1.0. The number of ID training and OOD test domains as well as their data volume are also identical to setups in Appendix D. The size of 3D spatial domain is $32\times 32\times 32$, and the past 10-step velocity field sequences are provided to forecast the future 11-step states. FNO3d is utilized as the backbone of operator experts for iMOOE3d. We present the 3D OOD dynamics forecasting results in Table 16. We observe that iMOOE can attain 17.36% and 39.26% decrease on nMSE and fRMSE compared to previous 3D neural operators. Such results further validate the effectiveness of proposed physics-guided PDE invariance learning to more complex 3D PDE dynamics.
F.2.3 Application to Real-world Time Series Prediction
@@
F.2.4 Application to PDE dynamics on irregular spatial domain
We leverage the Airfoil forecasting benchmark on non-uniform grid proposed in OFormer (Li et al., 2023b), where the underlying grid is divided by highly irregular triangular meshes. This Airfoil dataset contains 1000 training and 100 testing sequences with different inflow speed (Mach number) and angles of attack. The objective is to jointly predict future 22 steps velocity, density and pressure using past 4 steps states. One of the key advantages of iMOOE is the elegant compatibility across various neural operators. Hence, iMOOE can integrate neural operators that are able to handle irregular spatial domains or unstructured meshes as well, such as OFormer (Li et al., 2023b), Geo-FNO (Li et al., 2023c) and VCNeF (Hagnberger et al., 2024). We can easily enable iMOOE on irregular geometries by two tiny adaptations: i) The prior spatial derivatives on irregular grid are calculated by the finite element method. ii) The frequency enrichment loss relying on fast fourier transform on uniform grid is discarded. In Table 17, we present that when equipped with the proposed iMOOE method, three neural operators can exhibit better OOD forecasting accuracy on flows around Airfoil. We randomly showcase a OOD flow velocity forecast in Fig. 15, 16.
Table 17: OOD forecasting results on irregular spatial domains.
Operators Variants nMSE RMSE
VCNeF Naive 5.45e-2 2.22e-1
+iMOOE 5.08e-2 1.98e-1
Geo-FNO Naive 5.15e-2 2.16e-1
+iMOOE 4.35e-2 1.88e-1
OFormer Naive 4.95e-2 2.04e-1
+iMOOE 4.14e-2 1.79e-1
F.3 More Results on Sensitivity Analysis
By hyperparameter tuning, we empirically find the fixed setting $\lambda_{pred}=1.0,\lambda_{inv}=0.001,\lambda_{freq}=0.1,\lambda_{mask}=0.001$ can perform well on both simulated and real-world physical dynamics data. As the effect of $\mathcal{L}_{mask}$ has been discussed in Appendix F.1.2, we investigate iMOOE’s sensitivity to different $\lambda_{inv}$ and $\lambda_{freq}$ values. OOD results in Table 18 reflect that we should assign moderate values to $\lambda_{inv}$ and $\lambda_{freq}$ for satisfactory outcomes.
@@ F.3 More Results on Sensitivity Analysis
F.4 Further comparison with Meta-learning-based Methods
@@
F.5 Empirical Upper Bound of Zero-shot OOD Performance
@@
F.6 More Results on ID-OOD Correlations in Fig. 1(c)
@@
F.7 Analysis on Training Data Properties
@@
For the second question, we keep the total number of training samples at 1,024 and alter the number of training environments from 4 to 512. The training data quantity in each environment is equal. We depict the distribution of ID and OOD results of each test sample in Fig. 6(b), 7(b). Overall, with diverse training domains, i.e. when the number of training environments is up to 32, ID/OOD results of each test trajectory can disperse more compactly. In other words, the variance across test domains is smaller, and the average ID/OOD fRMSE is much lower. This reveals that when data budget is limited, better data diversity can avoid overfitting to limited training domains, and aid to find the fundamental PDE invariance principle by equalizing the risks across more diverse training environments. This is coherent with the key claim in foundational invariant learning literature (Arjovsky et al., 2019): with a sufficiently large number of diverse training environments, invariant risk minimization would elicit the invariant predictor.
@@ F.9 Implementation Details
F.10 Visualization on OOD Forecasting Results of iMOOE
In Fig. 8 to 14, we visualize the forecasting outcomes of iMOOE on representative OOD physical environments of the five PDE dynamical systems.
@@ F.10 Visualization on OOD Forecasting Results of iMOOE
Figure 15: OOD forecast showcase on the x-axis velocity around Airfoil with unseen conditions.
Figure 16: OOD forecast showcase on the y-axis velocity around Airfoil with unseen conditions.